# Multipoint distributions of the KPZ fixed point with compactly supported initial conditions

### v3

@@ Abstract
In this paper, we obtain the first formula for the space-time joint distributions of the KPZ fixed point with general initial conditions of compact support. The formula is obtained through taking $1:2:3$ KPZ scaling limit of the multipoint distribution formulas for the totally asymmetric simple exclusion process (TASEP). A key ingredient is a probabilistic representation, inspired by [MQR21], of the kernel encoding the initial condition for TASEP, which was first defined through an implicit characterization in [Liu22]. Moreover, we also verify that the equal time version of our formula matches the path integral formula in [MQR21] for the KPZ fixed point.
@@ 1.1 Background
The Kardar–Parisi–Zhang (KPZ) universality class [KPZ86] contains a broad family of random growth models in $(1+1)$-dimensions, including models from directed polymers, interacting particle systems, stochastic partial differential equations, etc. In the past four decades, the KPZ universality class has become a central subject of study in probability theory, statistical mechanics, and mathematical physics. For a more thorough introduction, we refer to the surveys [Cor12, Qua12, Zyg22] and the references therein.
@@ 1.2 Main results
Similar as in the narrow wedge case [JR21, Liu22], our multipoint formula for the KPZ fixed point with a general initial condition has the form of contour integrals of a Fredholm determinant. The Fredholm determinant $\mathbf{D}_{\mathfrak{h}}$ has a block diagonal kernel acting on nested Airy-type contours. The dependency on the initial condition is only through the top-left corner of the kernel, characterized by a function $\chi_{\mathfrak{h}}(\eta,\xi)$ defined on certain Airy contours, see Section 2.1.1 for its definition. For the narrow wedge case, our formula matches with the one in [Liu22].
@@ 1.3 Outline of the proof and some discussions
Theorem 1.2 is proved by taking a $1:2:3$ scaling limit of the corresponding multipoint distribution formulas of the totally asymmetric simple exclusion process (TASEP). The starting point is an algebraic formula obtained in [Liu22, Theorem 2.1] for the multipoint (space-time) distribution of TASEP starting from any right-finite initial condition. The dependency of the TASEP formula on the initial condition is encoded in a function $\mathrm{ch}_{Y}(v,u)$, which is characterized by an implicit reproducing-type property, see Definition 3.1. For the step and (pseudo) flat initial condition, an explicit form of $\mathrm{ch}_{Y}(v,u)$ in terms of symmetric functions was obtained in [Liu22] and is suitable for asymptotics, thus leading to the corresponding multipoint formula for the KPZ fixed point starting from the narrow wedge and flat initial conditions after taking limits.
@@
A key ingredient of this paper is an explicit probabilistic representation of the function $\mathrm{ch}_{Y}(v,u)$, through a hitting expectation with respect to geometric random walks, see Theorem 3.4. The probabilistic representation is suitable for asymptotic analysis and leads to the Brownian hitting representation in the limit. For technical reasons, we first take the limit of the TASEP formula under the assumption that the KPZ fixed point starts with initial conditions consisting of finitely many narrow wedges, and then extend the formula to compactly supported initial condition at the level of the KPZ fixed point, using a density argument and the continuity of the law of the KPZ fixed point with respect to initial conditions.
@@
Finally, we comment on our assumptions on the initial conditions. It would be desirable if one can get a formula that works for all initial conditions $\mathfrak{h}\in\mathrm{UC}$, in particular, the flat initial condition $\mathfrak{h}\equiv 0$. The reason we choose to restrict to the subspace $\mathrm{UC}_{c}$ is not merely a technical issue. There are genuine structural difficulties in this generality: the characteristic function $\chi_{\mathfrak{h}}(\eta,\xi)$ of the initial condition (see Definition 2.2) may not be well-defined pointwisely in general. Indeed, for the flat initial condition $\mathfrak{h}\equiv 0$, one can show from our formula that $\chi_{\mathfrak{h}}(\eta,\xi)$ is the dirac delta kernel $\delta_{\eta=-\xi}$, when properly interpreted. This suggests that in general, one should conjugate our formula to real spaces so that the limit when the support goes to infinity exists, at the level of operators acting on real spaces. We choose to stick to contour integral/Fourier type kernels, as it makes the algebraic structure of the multi-time formula much nicer and the analogy between narrow wedge and general initial conditions much more transparent. We leave it as a future project to extend our formula to any $\mathfrak{h}\in\mathrm{UC}$, with a kernel acting on real spaces.
@@ Acknowledgements
YL acknowledges the support of a Junior Fellowship at Institut Mittag-Leffler in Djursholm, Sweden, during the Fall semester of 2024, funded by the Swedish Research Council under grant no. 2021-06594. The work of ZL is partially supported by NSF DMS-2246683. The authors thank Konstantin Matetski, Li-Cheng Tsai, Xuan Wu, and Nikolaos Zygouras for their comments and suggestions.
@@ 2.1.1 The characteristic function $\chi_{\mathfrak{h}}$
For $\mathfrak{h}\in\mathrm{UC}_{c}$, the function $\chi_{\mathfrak{h}}(\eta,\xi)$ is well-defined and analytic for $\zeta\in\mathbb{C}_{\mathrm{L}}$ and $\eta\in\mathbb{C}_{\mathrm{R}}$. Moreover, if $\mathfrak{h}(\alpha)=$-$\infty$ for $|\alpha|>L$, we have
$|\chi_{\mathfrak{h}}(\eta,\xi)|\leq CL^{\frac{3}{2}}\mathrm{e}^{\mathrm{Re}(\eta-\xi)\cdot\max_{\alpha\in\mathbb{R}}\mathfrak{h}(\alpha)}\cdot\left(\mathrm{e}^{cL(|\xi|^{2}+|\eta|^{2})}+\frac{1}{\mathrm{Re}(\eta-\xi)}\right),$ (2.10)
for some constants $c,C>0$ depending only on $\mathfrak{h}$.
@@
We verify the statement for the first term appearing on the right-hand side of (2.9), the other two terms can be treated similarly. For the well-definedness it suffices to show that the integral over the $s$-variable converges absolutely. Take $L>0$ large enough such that $\mathrm{supp}(\mathfrak{h})\subset[-L,L]$. We split the integral into two parts depending on whether $s\leq\mathfrak{h}(0)$ or not. For $s\leq\mathfrak{h}(0)$ clearly we have $\bm{\tau}_{+}=0$ and $\mathbf{B}(\bm{\tau}_{+})=s$, hence
$\int_{-\infty}^{\mathfrak{h}(0)}\mathop{}\!\mathrm{d}s\,\mathrm{e}^{s\eta}\cdot\mathbb{E}_{\mathbf{B}(0)=s}\left[\exp\left(-\bm{\tau}_{+}\xi^{2}-\mathbf{B}(\bm{\tau}_{+})\xi\right)\mathbf{1}_{\bm{\tau}_{+}<\infty}\right]=\int_{-\infty}^{\mathfrak{h}(0)}\mathop{}\!\mathrm{d}s\,\mathrm{e}^{-s\xi+s\eta}=\frac{\mathrm{e}^{\mathfrak{h}(0)(\eta-\xi)}}{\eta-\xi},$ (2.11)
with the integral converging absolutely since $\mathrm{Re}(\xi)<\mathrm{Re}(\eta)$ by our assumption. For the other part, note that $\mathfrak{h}\in\mathrm{UC}_{c}$ is bounded above, namely, $\mathfrak{h}(\alpha)\leq\beta$ for some $\beta\in\mathbb{R}$. Define a new stopping time
$\bm{\sigma}_{+}:=\inf\{\alpha\geq 0:\mathbf{B}(\alpha)\leq\beta\},$
then clearly we have $\bm{\sigma}_{+}\leq\bm{\tau}_{+}$. Hence $\mathbb{P}_{\mathbf{B}(0)=s}(\bm{\tau}_{+}<+\infty)\leq\mathbb{P}_{\mathbf{B}(0)=s}(\bm{\sigma}_{+}<+\infty)$. Thus,
$\mathbb{E}_{\mathbf{B}(0)=s}\left[\left|\exp\left(-\bm{\tau}_{+}\xi^{2}-\mathbf{B}(\bm{\tau}_{+})\xi\right)\mathbf{1}_{\bm{\tau}_{+}<\infty}\right|\right]$ (2.12)
$\leq\mathbb{E}_{\mathbf{B}(0)=s}\left[\exp\left(\bm{\tau}_{+}\cdot\mathrm{Re}(-\xi^{2})-\beta\cdot\mathrm{Re}(\xi)\right)\mathbf{1}_{\bm{\tau}_{+}<\infty}\right]$
@@
$\leq\exp(L|\xi|^{2}-\beta\mathrm{Re}(\xi))\cdot\mathbb{E}_{\mathbf{B}(0)=s}\left[\mathbf{1}_{\bm{\tau}_{+}<\infty}\right]\leq\exp(L|\xi|^{2}-\beta\mathrm{Re}(\xi))\cdot\mathbb{E}_{\mathbf{B}(0)=s}\left[\mathbf{1}_{\bf{\bm{\sigma}}_{+}<\infty}\right].$
The density of $\bm{\sigma}_{+}$ can be computed using the reflection principle, see, e.g., [Dur19, (7.4.6)]. We have
$\mathbb{P}_{\mathbf{B}(0)=s}(\bm{\sigma}_{+}\in\mathop{}\!\mathrm{d}T)=\frac{|s-\beta|}{\sqrt{4\pi T^{3}}}\mathrm{e}^{-\frac{(s-\beta)^{2}}{4T}}\,\mathop{}\!\mathrm{d}T,$ (2.13)
whenever $s>\beta$. Thus for $\mathfrak{h}$ supported on [L,L][-L,L], we have
$\int_{\mathfrak{h}(0)}^{\infty}\mathop{}\!\mathrm{d}s\,\left|\mathrm{e}^{s\eta}\right|\cdot\mathbb{E}_{\mathbf{B}(0)=s}\left[\left|\exp\left(-\bm{\tau}_{+}\xi^{2}-\mathbf{B}(\bm{\tau}_{+})\xi\right)\right|\mathbf{1}_{\bm{\tau}_{+}<\infty}\right]$ (2.14)
$\leq\mathrm{e}^{L|\xi|^{2}}\int_{\mathfrak{h}(0)}^{\beta+1}\mathop{}\!\mathrm{d}s\,\mathrm{e}^{s\mathrm{Re}(\eta)-\beta\mathrm{Re}(\xi)}+\mathrm{e}^{L|\xi|^{2}}\int_{\beta+1}^{\infty}\mathop{}\!\mathrm{d}s\int_{0}^{L}\mathop{}\!\mathrm{d}T\,\mathrm{e}^{s\mathrm{Re}(\eta)}\cdot\frac{|s-\beta|}{\sqrt{4\pi T^{3}}}\mathrm{e}^{-\frac{(s-\beta)^{2}}{4T}-\beta\cdot\mathrm{Re}(\xi)}$
$\leq C_{1}\mathrm{e}^{L|\xi|^{2}+\mathrm{Re}(\eta)}\mathrm{e}^{\beta\cdot\mathrm{Re}(\eta-\xi)}+C_{2}\mathrm{e}^{L|\xi|^{2}}\int_{\beta+1}^{\infty}\mathop{}\!\mathrm{d}s\int_{0}^{L}\mathop{}\!\mathrm{d}T\ \mathrm{e}^{s\mathrm{Re}(\eta)-\frac{(s-\beta)^{2}}{8T}-\beta\cdot\mathrm{Re}(\xi)},$
where $C_{1}$ and $C_{2}$ are two constants that depend on $\beta$, and in the last step we used the elementary inequality
$\frac{|s-\beta|}{\sqrt{4\pi T^{3}}}\mathrm{e}^{-\frac{(s-\beta)^{2}}{4T}}=\frac{|s-\beta|}{\sqrt{4\pi T^{3}}}\mathrm{e}^{-\frac{(s-\beta)^{2}}{8T}}\cdot\mathrm{e}^{-\frac{(s-\beta)^{2}}{8T}}\leq\max_{T\in[0,\infty)}\left\{\frac{|s-\beta|}{\sqrt{4\pi T^{3}}}\mathrm{e}^{-\frac{(s-\beta)^{2}}{8T}}\right\}\cdot\mathrm{e}^{-\frac{(s-\beta)^{2}}{8T}}\leq\frac{C}{(s-\beta)^{2}}\cdot\mathrm{e}^{-\frac{(s-\beta)^{2}}{8T}}.$
Now by bounding $\mathrm{e}^{-\frac{(s-\beta)^{2}}{8T}}$ by $\mathrm{e}^{-\frac{(s-\beta)^{2}}{8L}}$ for $T\in[0,L]$, we have
$\int_{\beta+1}^{\infty}\mathop{}\!\mathrm{d}s\int_{0}^{L}\mathop{}\!\mathrm{d}T\ \mathrm{e}^{s\mathrm{Re}(\eta)-\frac{(s-\beta)^{2}}{8T}-\beta\cdot\mathrm{Re}(\xi)}\leq L\mathrm{e}^{-\beta\cdot\mathrm{Re}(\xi)}\int_{\beta+1}^{\infty}\mathop{}\!\mathrm{d}s\ \mathrm{e}^{s\mathrm{Re}(\eta)-\frac{(s-\beta)^{2}}{8L}}\leq CL^{\frac{3}{2}}\mathrm{e}^{\beta\cdot\mathrm{Re}(\eta-\xi)+2L|\eta|^{2}}.$
Thus, we conclude that
@@
$\begin{split}&\left|\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathrm{e}^{s\eta}\cdot\mathbb{E}_{\mathbf{B}(0)=s}\left[\exp\left(-\bm{\tau}_{+}\xi^{2}-\mathbf{B}(\bm{\tau}_{+})\xi\right)\mathbf{1}_{\bm{\tau}_{+}<\infty}\right]\right|\\ &\leq CL^{\frac{3}{2}}\mathrm{e}^{\mathrm{Re}(\eta-\xi)\cdot\max_{\alpha\in\mathbb{R}}\mathfrak{h}(\alpha)}\cdot\left(\mathrm{e}^{cL(|\xi|^{2}+|\eta|^{2})}+\frac{1}{\mathrm{Re}(\eta-\xi)}\right).\end{split}$
The estimates for the other two parts are similar. The analyticity in $\xi$ and $\eta$ is an easy consequence of the above bound and the dominated convergence theorem.
@@
For convenience, we were using the origin 0 as the starting point of the Brownian motions in the hitting formula. The next proposition shows that this is not necessary, and one can start with any point on the real line $\mathbb{R}$ and get the same characteristic function $\chi_{\mathfrak{h}}$.

### v4

@@ Abstract
In this paper, we obtain the first formula for the space-time joint distributions of the KPZ fixed point with general initial conditions of compact support. The formula is obtained through taking $1:2:3$ KPZ scaling limit of the multipoint distribution formulas for the totally asymmetric simple exclusion process (TASEP). A key ingredient is a probabilistic representation, inspired by [MQR21], of the kernel encoding the initial condition for TASEP, which was first defined through an implicit characterization in [Liu22]. Moreover, we also verify that the equal time version of our formula matches the path integral formula in [MQR21] for the KPZ fixed point when the initial condition is of compact support.
@@ 1.1 Background
The Kardar–Parisi–Zhang (KPZ) universality class [KPZ86] contains a broad family of random growth models in $(1+1)$-dimensions, including models from directed polymers [Sep12, COSZ14], interacting particle systems [Joh00], stochastic partial differential equations [KPZ86, Hai13], etc. In the past four decades, the KPZ universality class has become a central subject of study in probability theory, statistical mechanics, and mathematical physics. For a more thorough introduction, we refer to the surveys [Cor12, Qua12, Zyg22] and the references therein.
@@ 1.2 Main results
Similar as in the narrow wedge case [JR21, Liu22], our multipoint formula for the KPZ fixed point with a general initial condition has the form of contour integrals of a Fredholm determinant. The Fredholm determinant $\mathbf{D}_{\mathfrak{h}}$ has a block diagonal kernel acting on nested Airy-type contours. The dependency on the initial condition is only through the top-left corner of the kernel, characterized by a function $\chi_{\mathfrak{h}}(\eta,\xi)$ defined on certain Airy contours, see Section 2.1.1 for its definition. For the narrow wedge case, our formula recovers the one in [Liu22].
@@ 1.3 Outline of the proof and some discussions
Theorem 1.2 is proved by taking a $1:2:3$ scaling limit of the corresponding multipoint distribution formulas of the totally asymmetric simple exclusion process (TASEP). The starting point is an algebraic formula obtained in [Liu22, Theorem 2.1] for the multipoint (space-time) distribution of TASEP starting from any right-finite initial condition. The dependency of the TASEP formula on the initial condition is encoded in a function $\mathrm{ch}_{Y}(v,u)$, which is characterized by an implicit reproducing-type property, see Definition 3.1. An explicit form of $\mathrm{ch}_{Y}(v,u)$ in terms of symmetric functions was also obtained in [Liu22], which is suitable for asymptotics for the step and (pseudo) flat initial conditions. Thus it led to the corresponding multipoint formula in [Liu22] for the KPZ fixed point starting from the narrow wedge and flat initial conditions after taking limits.
@@
A key ingredient of this paper is an explicit probabilistic representation of the function $\mathrm{ch}_{Y}(v,u)$, through a hitting expectation with respect to geometric random walks, see Theorem 3.4. The probabilistic representation is suitable for asymptotic analysis for more general initial conditions and leads to the Brownian hitting representation in the limit. For technical reasons, we first take the limit of the TASEP formula under the assumption that the KPZ fixed point starts with initial conditions consisting of finitely many narrow wedges, and then extend the formula to compactly supported initial condition at the level of the KPZ fixed point, using a density argument and the continuity of the law of the KPZ fixed point with respect to initial conditions.
@@
Finally, we comment on our assumptions on the initial conditions. It would be desirable if one can get a formula that works for all initial conditions $\mathfrak{h}\in\mathrm{UC}$, in particular, the flat initial condition $\mathfrak{h}\equiv 0$. The reason we choose to restrict to the subspace $\mathrm{UC}_{c}$ is not merely a technical issue. There are genuine structural difficulties in this generality: the characteristic function $\chi_{\mathfrak{h}}(\eta,\xi)$ of the initial condition (see Definition 2.2) may not be well-defined pointwisely in general. It might be possible to define $\chi_{\mathfrak{h}}(\eta,\xi)$ in a proper sense case by case when $\mathfrak{h}$ is not compactly supported. Indeed, for the flat initial condition $\mathfrak{h}\equiv 0$, one can show from our formula that $\chi_{\mathfrak{h}}(\eta,\xi)$ is the dirac delta integral kernel $\delta_{\eta=-\xi}$ [Liu22]. However, our current strategy requires to control the rate of the growth of the kernel, where compactness of the initial condition is needed. It might be possible that one needs to interpret $\chi_{\mathfrak{h}}(\eta,\xi)$ in the sense of distribution instead of a function. It might also be possible to conjugate our formula to real spaces so that the limit when the support goes to infinity exists, at the level of operators acting on real spaces. We leave it as a future project to extend our formula to any $\mathfrak{h}\in\mathrm{UC}$, with a proper way to understand $\chi_{\mathfrak{h}}(\eta,\xi)$.
@@ Acknowledgements
YL acknowledges the support of a Junior Fellowship at Institut Mittag-Leffler in Djursholm, Sweden, during the Fall semester of 2024, funded by the Swedish Research Council under grant no. 2021-06594. The work of ZL is partially supported by NSF DMS-2246683. The authors thank Konstantin Matetski, Daniel Remenik, Tejaswi Tripathi, Li-Cheng Tsai, Xuan Wu, and Nikolaos Zygouras for their comments and suggestions.
@@ 2.1.1 The characteristic function $\chi_{\mathfrak{h}}$
We remark that when $\mathfrak{h}(\alpha)=-\infty\mathbf{1}_{\alpha\neq 0}$ is the narrow wedge initial condition, the above characteristic function can be evaluated as $\chi_{\mathfrak{h}}(\eta,\xi)=\frac{1}{\eta-\xi}$, which matches the corresponding function in [Liu22] for the narrow wedge initial condition.
@@
For $\mathfrak{h}\in\mathrm{UC}_{c}$, the function $\chi_{\mathfrak{h}}(\eta,\xi)$ is well-defined and analytic for $\zeta\in\mathbb{C}_{\mathrm{L}}$ and $\eta\in\mathbb{C}_{\mathrm{R}}$. Moreover, for any $L>0$ and $\beta\in\mathbb{R}$ such that $\mathrm{supp}(\mathfrak{h})\subset[-L,L]$ and $\max_{\alpha\in\mathbb{R}}\mathfrak{h}(\alpha)\leq\beta$, we have
$|\chi_{\mathfrak{h}}(\eta,\xi)|\leq\mathrm{e}^{(\beta+1)\mathrm{Re}(\eta-\xi)+2L(|\xi|^{2}+|\eta|^{2})}\cdot\left(\frac{2}{\mathrm{Re}(\eta)}+\frac{2}{\mathrm{Re}(-\xi)}+8L+\frac{2^{5/2}L^{3/2}}{\sqrt{\pi}}\right).$ (2.10)
@@
Assume $\mathrm{supp}(\mathfrak{h})\subset[-L,L]$ and $\max_{\alpha\in\mathbb{R}}\mathfrak{h}(\alpha)\leq\beta$. We will show that the integrals on the right-hand side of (2.9) are absolutely convergent and uniformly bounded by the right-hand side of (2.10). Therefore $\chi_{\mathfrak{h}}$ is well defined and analytic on $\mathbb{C}_{\mathrm{R}}\times\mathbb{C}_{\mathrm{L}}$.
We consider the first integral on the right-hand side of (2.9). Note that $\mathbf{1}_{\tau_{+}<\infty}=\mathbf{1}_{\tau_{+}\leq L}$, and $\mathrm{Re}(-\bm{\tau}_{+}\xi^{2}-\mathbf{B}(\bm{\tau}_{+})\xi)$ is bounded by $L|\xi^{2}|+\beta\mathrm{Re}(-\xi)$ when $\tau_{+}\leq L$. These facts imply
@@
$\begin{split}&\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,|\mathrm{e}^{+s\eta}|\cdot\mathbb{E}_{\mathbf{B}(0)=s}\left[|\exp\left(-\bm{\tau}_{+}\xi^{2}-\mathbf{B}(\bm{\tau}_{+})\xi\right)|\mathbf{1}_{\bm{\tau}_{+}<\infty}\right]\\ &\leq\mathrm{e}^{L|\xi^{2}|-\beta\mathrm{Re}(\xi)}\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathrm{e}^{s\mathrm{Re}(\eta)}\cdot\mathbb{P}_{\mathbf{B}(0)=s}(\bm{\tau}_{+}\leq L)\\ &\leq\mathrm{e}^{L|\xi^{2}|-\beta\mathrm{Re}(\xi)}\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathrm{e}^{s\mathrm{Re}(\eta)}\cdot\mathbb{P}_{\mathbf{B}(0)=s}(\bm{\sigma}_{+}\leq L),\end{split}$ (2.11)
where we used the following fact that $\bm{\tau}_{+}\geq\bm{\sigma}_{+}$ with $\bm{\sigma}_{+}$ being the following new stopping time
$\bm{\sigma}_{+}:=\inf\{\alpha\geq 0:\mathbf{B}(\alpha)\leq\beta\}.$
Note that the distribution of $\bm{\sigma}_{+}$ can be computed using the reflection principle, see, e.g., [Dur19, (7.4.4)]. When $s\geq\beta+1$, we can estimate
@@
$\begin{split}\mathbb{P}_{\mathbf{B}(0)=s}(\bm{\sigma}_{+}\leq L)&=2\mathbb{P}_{\mathbf{B}(0)=s}\left(\mathbf{B}(L)\leq\beta\right)\\ &=\frac{1}{\sqrt{\pi L}}\int_{-\infty}^{\beta}\mathrm{e}^{-\frac{(y-s)^{2}}{4L}}\mathop{}\!\mathrm{d}y=\frac{1}{\sqrt{\pi L}}\mathrm{e}^{-\frac{(\beta-s)^{2}}{4L}}\int_{-\infty}^{0}e^{-\frac{(\beta-s)z}{2L}-\frac{z^{2}}{4L}}\mathop{}\!\mathrm{d}z\\ &\leq\frac{1}{\sqrt{\pi L}}\mathrm{e}^{-\frac{(\beta-s)^{2}}{4L}}\int_{-\infty}^{0}e^{-\frac{(\beta-s)z}{2L}}\mathop{}\!\mathrm{d}z=\frac{2\sqrt{L}}{\sqrt{\pi}(s-\beta)}\mathrm{e}^{-\frac{(\beta-s)^{2}}{4L}}\\ &\leq\frac{2\sqrt{L}}{\sqrt{\pi}}\mathrm{e}^{-\frac{(\beta-s)^{2}}{4L}}.\end{split}$ (2.12)
Therefore,
$\begin{split}\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathrm{e}^{s\mathrm{Re}(\eta)}\cdot\mathbb{P}_{\mathbf{B}(0)=s}(\bm{\sigma}_{+}\leq L)&\leq\int_{-\infty}^{\beta+1}\mathop{}\!\mathrm{d}s\,\mathrm{e}^{s\mathrm{Re}(\eta)}+\frac{2\sqrt{L}}{\sqrt{\pi}}\int_{\beta+1}^{\infty}\mathop{}\!\mathrm{d}s\,\mathrm{e}^{s\mathrm{Re}(\eta)-\frac{(\beta-s)^{2}}{4L}}\\ &\leq\frac{\mathrm{e}^{(\beta+1)\mathrm{\mathrm{Re}}(\eta)}}{\mathrm{Re}(\eta)}+4L\mathrm{e}^{L(\mathrm{Re}(\eta))^{2}+\beta\mathrm{Re}(\eta)}.\end{split}$ (2.13)
By inserting this bound to (2.11) and noting that $(\mathrm{Re}(\eta))^{2}\leq|\eta^{2}|$, we obtain
$\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,|\mathrm{e}^{+s\eta}|\cdot\mathbb{E}_{\mathbf{B}(0)=s}\left[|\exp\left(-\bm{\tau}_{+}\xi^{2}-\mathbf{B}(\bm{\tau}_{+})\xi\right)|\mathbf{1}_{\bm{\tau}_{+}<\infty}\right]\leq\mathrm{e}^{(\beta+1)\mathrm{Re}(\eta-\xi)+L(|\xi|^{2}+|\eta|^{2})}\cdot\left(4L+\frac{1}{\mathrm{Re}(\eta)}\right).$ (2.14)
The second integral on the right-hand side of (2.9) can be handled similarly. We get
$\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,|\mathrm{e}^{-s\xi}|\cdot\mathbb{E}_{\mathbf{B}(0)=s}\left[|\exp\left(\bm{\tau}_{-}\eta^{2}+\mathbf{B}(\bm{\tau}_{-})\eta\right)|\mathbf{1}_{\bm{\tau}_{-}>-\infty}\right]\leq\mathrm{e}^{(\beta+1)\mathrm{Re}(\eta-\xi)+L(|\xi|^{2}+|\eta|^{2})}\cdot\left(4L+\frac{1}{\mathrm{Re}(-\xi)}\right).$ (2.15)
For the last integral, we need to slightly modify the estimate. We need a different estimate of the integral when $s\leq\beta+1$. We still have $\tau_{+}\leq L$ when $\tau_{+}<\infty$, and $\tau_{-}\geq-L$ when $\tau_{-}>$-$\infty$. Also note that $\mathbf{B}(\tau_{+})\leq\max_{t\in[0,L]}\mathbf{B}(t)$ and $\mathbf{B}(\tau_{-})\leq\max_{t\in[-L,0]}\mathbf{B}(t)$. Thus we have
$\begin{split}&\int_{-\infty}^{\beta+1}\mathop{}\!\mathrm{d}s\,\mathbb{E}_{\mathbf{B}(0)=s}\left[|\exp\left(-\bm{\tau}_{+}\xi^{2}-\mathbf{B}(\bm{\tau}_{+})\xi\right)|\mathbf{1}_{\bm{\tau}_{+}<\infty}\right]\cdot\mathbb{E}_{\mathbf{B}(0)=s}\left[|\exp\left(\bm{\tau}_{-}\eta^{2}+\mathbf{B}(\bm{\tau}_{-})\eta\right)|\mathbf{1}_{\bm{\tau}_{-}>-\infty}\right]\\ &\leq\mathrm{e}^{L(|\xi^{2}|+|\eta^{2}|)}\int_{-\infty}^{\beta+1}\mathop{}\!\mathrm{d}s\,\mathbb{E}_{\mathbf{B}(0)=s}\left[\mathrm{e}^{\mathrm{Re}(-\xi)\max_{t\in[0,L]}\mathbf{B}(t)}\right]\cdot\mathbb{E}_{\mathbf{B}(0)=s}\left[\mathrm{e}^{\mathrm{Re}(\eta)\max_{t\in[-L,0]}\mathbf{B}(t)}\right]\\ &=\mathrm{e}^{L(|\xi^{2}|+|\eta^{2}|)}\int_{-\infty}^{\beta+1}\mathop{}\!\mathrm{d}s\,\mathrm{e}^{s\mathrm{Re}(\eta-\xi)}\cdot\mathbb{E}_{\mathbf{B}(0)=0}\left[\mathrm{e}^{\mathrm{Re}(-\xi)\max_{t\in[0,L]}\mathbf{B}(t)}\right]\cdot\mathbb{E}_{\mathbf{B}(0)=0}\left[\mathrm{e}^{\mathrm{Re}(\eta)\max_{t\in[-L,0]}\mathbf{B}(t)}\right]\\ &=\mathrm{e}^{L(|\xi^{2}|+|\eta^{2}|)}\cdot\frac{\mathrm{e}^{(\beta+1)\mathrm{Re}(\eta-\xi)}}{\mathrm{Re}(\eta-\xi)}\cdot\left(\frac{1}{\sqrt{\pi L}}\int_{0}^{\infty}\mathrm{e}^{-\frac{x^{2}}{4L}+\mathrm{Re}{(-\xi)}x}\mathop{}\!\mathrm{d}x\right)\cdot\left(\frac{1}{\sqrt{\pi L}}\int_{0}^{\infty}\mathrm{e}^{-\frac{x^{2}}{4L}+\mathrm{Re}(\eta)x}\mathop{}\!\mathrm{d}x\right)\\ &\leq\mathrm{e}^{L(|\xi^{2}|+|\eta^{2}|)}\cdot\frac{\mathrm{e}^{(\beta+1)\mathrm{Re}(\eta-\xi)}}{\mathrm{Re}(\eta-\xi)}\cdot\mathrm{e}^{L(\mathrm{Re}{(-\xi}))^{2}+L(\mathrm{Re}(\eta))^{2}}\\ &\leq\mathrm{e}^{2L(|\xi^{2}|+|\eta^{2}|)+(\beta+1)\mathrm{Re}(\eta-\xi)}\cdot\left(\frac{1}{\mathrm{Re}(-\xi)}+\frac{1}{\mathrm{Re}(\eta)}\right),\end{split}$ (2.16)
where we used the fact that $\max_{t\in[0,L]}\mathbf{B}(t)$ and $\max_{t\in[-L,0]}\mathbf{B}(t)$ have the same distribution as $|\mathbf{B}(L)|$ when $\mathbf{B}(0)=0$. When $s>\beta+1$, we use the following bound which is similar to (2.11) and (2.13),
$\begin{split}&\int_{\beta+1}^{\infty}\mathop{}\!\mathrm{d}s\,\mathbb{E}_{\mathbf{B}(0)=s}\left[|\exp\left(-\bm{\tau}_{+}\xi^{2}-\mathbf{B}(\bm{\tau}_{+})\xi\right)|\mathbf{1}_{\bm{\tau}_{+}<\infty}\right]\cdot\mathbb{E}_{\mathbf{B}(0)=s}\left[|\exp\left(\bm{\tau}_{-}\eta^{2}+\mathbf{B}(\bm{\tau}_{-})\eta\right)|\mathbf{1}_{\bm{\tau}_{-}>-\infty}\right]\\ &\leq\mathrm{e}^{\beta\mathrm{Re}(\eta-\xi)+L(|\xi^{2}|+|\eta|^{2})}\int_{\beta+1}^{\infty}\mathop{}\!\mathrm{d}s\,\mathbb{P}_{\mathbf{B}(0)=s}(\bm{\sigma}_{+}\leq L)\mathbb{P}_{\mathbf{B}(0)=s}(\bm{\sigma}_{-}\geq-L)\\ &\leq\mathrm{e}^{\beta\mathrm{Re}(\eta-\xi)+L(|\xi^{2}|+|\eta|^{2})}\cdot\frac{4L}{\pi}\int_{\beta+1}^{\infty}\mathop{}\!\mathrm{d}s\,\mathrm{e}^{-\frac{(\beta-s)^{2}}{2L}}\\ &\leq\mathrm{e}^{\beta\mathrm{Re}(\eta-\xi)+L(|\xi^{2}|+|\eta|^{2})}\cdot\frac{2^{5/2}L^{3/2}}{\sqrt{\pi}},\end{split}$ (2.17)
where $\bm{\sigma}_{-}:=\sup\{\alpha\leq 0:\mathbf{B}(\alpha)\leq\beta\}$ is a stopping time which has the same law as $\bm{\sigma}_{+}$. By combining all the bounds (2.14), (2.15), (2.16) and (2.17), we prove the desired bound (2.10).
@@
For convenience, we were using the origin 0 as the starting point of the Brownian motions in the hitting expectation formula. The next proposition shows that this is not necessary, and one can start with any point on the real line $\mathbb{R}$ and get the same characteristic function $\chi_{\mathfrak{h}}$.